OpenStudy (anonymous):

Can somone please refresh my memory on how to put 6/200 into percents. I know the answer is 3% put i cannot remember for the life of me why! ^_^

OpenStudy (anonymous):

6/200 x 100

OpenStudy (anonymous):

Ok please explain a little more buddy? ^_^

OpenStudy (anonymous):

any fraction uve got u multiply it by 100 to get percentage...

OpenStudy (anonymous):

Ik that lol wow *blonde moment* my computer jumbled evertthing up >.< cuz im on my crapppy comp :P

OpenStudy (anonymous):

(6/200)%=6/2oo*100=6/2=3

OpenStudy (anonymous):

r u here

OpenStudy (anonymous):

there****

OpenStudy (anonymous):

yesh lol

OpenStudy (anonymous):

last time i m asking u:::: will u be my true and sincer frnd:... i need a shoulder to cry on....... my gf dumped me; i love her more thn any thng in the world and wat she did!!!!!! alasssssssssssssssssssssssssssssssssssssssssssssssss

OpenStudy (anonymous):

:o She did?!?!!? Ahwwwwwwwwwwwww :o thats rly terrible and sure i will be ur true friend of course!! :)

OpenStudy (anonymous):

thanks frnd

OpenStudy (anonymous):

weeeelccoomeeeee!!!! :D lol

OpenStudy (anonymous):

fb;;; i wana chat

OpenStudy (anonymous):

>.< i cant im not allowed to have one and i already got caught with one when i wasnt supposed to have one... :P soweeeee! >.<

OpenStudy (anonymous):

y............... why r u not allowed

OpenStudy (anonymous):

bc i got in trouble idk its a long story ><

OpenStudy (anonymous):

email??????????/

OpenStudy (anonymous):

lol i prolly wont answer it rite away but this is my RLY old one my mom didnt change the passsss its poisoniv1996@hotmail.com :)

OpenStudy (anonymous):

wana chat????

OpenStudy (anonymous):

r u there??????????????????????????????????????????????????????????